# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ A Macroscopic Thermodynamics
The total mean macroscopic EPR is $\langle\dot{\Sigma}\rangle=\langle\dot{\Sigma}^{\mathcal{D}}\rangle+\langle\dot{\Sigma}^{\mathcal{R}}\rangle+d_{t}S^{gb}$. It is decomposed into reactive mean macroscopic bulk EPR ($\langle\dot{\Sigma}^{\mathcal{R}}\rangle$), diffusive mean macroscopic bulk EPR ($\langle\dot{\Sigma}^{\mathcal{D}}\rangle$) and the boundary terms Gibbs EPR ($d_{t}S^{gb}$):
@@
Where $\mathbb{D}[\{\rho\}]$ denotes the path integral over the macrostate space and $\mathcal{P}[\{\rho\}]$ is the probability distribution for the macrostate. The mean EPR in eq. 33 has the form transition current multiplied by the transition affinity. The transition affinity is obtained using the macroscopic LDB eq. 27. The macroscopic Gibbs entropy is defined as $S^{gb}=-\int\mathbb{D}[\{\rho\}]\mathcal{P}[\{\rho\}]\ln{(\mathcal{P}[\{\rho\}])}$, which is related to the macroscopic state entropy $S^{state}$ is defined as [18, 20]:
@@ B Conservative and non-conservative decomposition of the total macroscopic EPR
$\langle\dot{\Sigma}\rangle$ is decomposed into four contributions analogous to eq. 13.
@@
The eq. 35 is the macroscopic second law of thermodynamics. $-d_{t}\langle E\rangle+d_{t}S^{gb}$ is the conservative EPR $\langle\dot{\Sigma}^{c}\rangle$, hence $\Sigma^{c}$ depends only on the initial and final state. The exact expression for $\langle\dot{\Sigma}^{ch}\rangle$ and $\langle\dot{\Sigma}^{sp}\rangle$ reads:
@@
The non-reciprocal EPR $\langle\dot{\Sigma}^{nr}\rangle$ consists of the reactive $\langle\dot{\Sigma}_{\mathcal{R}}^{nr}\rangle=\int_{\mathcal{V}}\sum_{\{\gamma\gamma^{\prime}\}}\langle J_{\gamma\gamma^{\prime}}\left(F^{nr}_{\gamma^{\prime}}-F^{nr}_{\gamma}\right)\rangle$ and diffusive EPR contributions $\langle\dot{\Sigma}^{nr}_{\mathcal{D}}\rangle=-\int_{\mathcal{V}}\sum_{i}\langle\vec{J}_{i}^{\vec{\mathcal{D}}}\cdot\nabla F^{nr}_{i}\rangle$, such that $\dot{\Sigma}^{nr}=\dot{\Sigma}_{\mathcal{R}}^{nr}+\dot{\Sigma}^{nr}_{\mathcal{D}}$. Using the EOM eq. 28 leads to $\dot{\Sigma}^{nr}=-\sum_{i}\int_{\mathcal{V}}F_{i}^{nr}\>\partial_{t}\rho_{i}$. Using $V_{ij}^{nr}=-V_{ji}^{nr}$ further leads to the simplified form:
@@
Here, $\omega_{ij}=\rho_{i}\partial_{t}\rho_{j}-\rho_{j}\partial_{t}\rho_{i}$ defines the macroscopic vorticity between $\rho_{i}$ and $\rho_{j}$. $\langle{\Sigma}^{nr}\rangle$ corresponds to the sustaining of vorticity currents between the macrostate generated by the macroscopic non-reciprocal forces. Importantly, the vorticity currents are defined in the macrostate space, in contrast to the dissipative currents defined in the transition space. Moreover, vorticity currents are defined using two macrostates, in contrast to chemical reaction networks that require a transition cycle of three states to formulate a dissipative current [1]. From a fundamental point of view, this reveals the sophisticated mechanism of non-equilibriumness in non-reciprocal systems, whose equivalent counterparts do not exist in known literature.
The dynamics of non-reciprocal systems exhibit dynamic phases due to $\mathcal{PT}$ symmetry breaking [54], for instance, travelling waves and temporal oscillations. Here, we focus on non-reciprocal phase transitions that exhibit a transition from a static to a dynamic phase. A static phase is defined as the $\mathcal{PT}$ symmetry-preserving phase, in contrast, the $\mathcal{PT}$ symmetry is broken in the dynamic phase. We consider a wave solution $\rho_{i}=\rho_{i}^{avg}+\phi_{i}^{ow}\cos{(\theta_{i}^{ow}-v^{ow}t)}+\phi_{i}^{tw}\cos{(x-v^{tw}t+\theta_{i}^{tw})}$. Here, $\rho_{i}^{avg}$ is the time and space integrated average value of $\rho_{i}$. Where, $w=\{tw,ow\}$ denotes the travelling wave and temporal oscillating wave respectively. The $\phi_{i}^{ow}$ and $\phi_{i}^{tw}$ characterize the wave amplitude. Similarly, $v_{i}^{ow}$ and $v_{i}^{tw}$ are the wave velocities. Integrating eq. 37 over space and oscillation time period $\tau^{ow}$, the non-reciprocal EP for the system reads:
@@
The eq. 38 quantifies the non-reciprocal EPR density over total volume and time-period. For $V_{ij}^{nr}\geq 0$, physically implies that $\rho_{i}$ repeals from $\rho_{j}$ and $\rho_{j}$ is attracted towards $\rho_{i}$, thus $\theta_{j}^{tw}\geq\theta_{i}^{tw}$ and $\theta_{j}^{ow}\geq\theta_{i}^{ow}$, which further implies $\langle{\Sigma}^{nr}\rangle\geq 0$. Importantly, the out-of-phase state $\theta_{j}^{tw}-\theta_{i}^{tw}=\pi/2$ maximizes $\langle{\Sigma}^{nr}\rangle$ and the in-phase state $\theta_{j}^{tw}=\theta_{i}^{tw}$ minimizes $\langle{\Sigma}^{nr}\rangle$. The in-phase to out-of-phase steady-state transition has been a key motif of non-reciprocal systems [54]. Hence, it signifies that the transition from the in-phase to out-of-phase steady-state is equivalent to switching the minimum to the maximum nonreciprocal EPR for steady-state selection. Moreover, $\langle\Sigma^{nr}\rangle$ for out-of-phase oscillation scales with observation time $\tau$. This signifies a constant thermodynamic dissipation cost required to sustain the vorticity currents between the macrostates. Physically, it connects the dynamical phase transition analogous to equilibrium phase transitions, where EPR for the non-equilibrium phase transition is analogous to the free energy for the equilibrium phase transition [135]. It is characterized by different scaling regimes for $\langle\Sigma^{nr}\rangle$. The dissipative nature of the vorticity currents is realized only in the dynamic phase.
The scaling of different contributions of $\Sigma$ with observation time $\tau$ is summarized in table 1. The self-propulsion and chemical driving EP follows scaling $O(\tau)$ for dissipative currents [88]. However, a discontinuity (or a kink) in self-propulsion or chemical driving EPR can be observed and associated with a dynamical phase transition [135, 136, 137, 138]. The non-reciprocal phase transitions are characterized by different scaling regimes for $\Sigma^{nr}$, which corresponds to sustaining temporal or spatial vorticity currents.
@@ C Orthogonal decomposition of the EPR
Using the definition eq. 39, $-d_{t}\langle{E}\rangle+d_{t}S^{gb}$ in eq. 35 is further decomposed to the orthogonal form [96]:
@@
$\psi_{E}$ is the macroscopic free energy defined as $\psi_{E}=-\log{(\mathcal{Z}^{E})}$. Here, $\mathcal{Z}^{E}=\int\mathbb{D}[\{\rho\}]e^{-E}$ is the partition function and the macroscopic Boltzmann distribution is ${\mathcal{P}^{E}[\{\rho\}]}=e^{-E\>+\Psi^{E}}$. $\langle S^{state}_{E}\rangle=-D^{KL}_{{E}}(\mathcal{P}[\{\rho\}])$ holds. Utilizing the symmetry of the non-conservative external driving forces $F_{\gamma\gamma^{\prime}}$ and $\vec{F}_{i}^{sp}$ namely the orthogonal decomposition of the transition affinities, eq. 36 is reduced to:
@@
Where $\vec{J}_{i}^{sp}=2d_{i}e^{\mu_{i}}\sinh{(\vec{F}_{i}^{sp}/2)}=T_{i}^{\perp}\sinh{(\vec{F}_{i}^{sp}/2)}$ is the anti-symmetric part of $\vec{J}_{i}^{\vec{\mathcal{D}}}$ under the adjoint transformation $\vec{F}_{i}^{sp}\to-\vec{F}_{i}^{sp}$. Similarly, $J_{\gamma\gamma^{\prime}}^{ch}=d_{\gamma\gamma^{\prime}}\left(e^{\mu_{\gamma}}+e^{\mu_{\gamma^{\prime}}}\right)\sinh{\left({F_{\gamma\gamma^{\prime}}^{ch}}/{2}\right)}=T^{\perp}_{\gamma\gamma^{\prime}}\sinh{\left({F_{\gamma\gamma^{\prime}}^{ch}}/{2}\right)}$ is the anti-symmetric part of $J_{\gamma\gamma^{\prime}}$ under the adjoint transformation $F_{\gamma\gamma^{\prime}}^{ch}\to-F_{\gamma\gamma^{\prime}}^{ch}$. Importantly, $T_{i}^{\perp}$ and $T^{\perp}_{\gamma\gamma^{\prime}}$ are the scaled (by $\Omega$) variance of currents (eq. 31) in the direction orthogonal to the external driving, which are obtained by plugging in $\vec{F}_{i}^{sp}=0$ and $F_{\gamma\gamma^{\prime}}^{ch}=0$ respectively. Thus in eq. 41, the non-conservative EPR depends on a non-linear function of the external driving force and the current variance in the orthogonal direction to the driving. Importantly, the $\langle\dot{\Sigma}^{ch}\rangle$ and $\langle\dot{\Sigma}^{sp}\rangle$ depend on $e^{\mu_{i}}$ and is propostional to the macrostate mobility.
@@
The eq. 40 formulates the orthogonal decomposition of $\langle\dot{\Sigma}\rangle$. It decomposes $\langle\dot{\Sigma}\rangle$ into four linearly independent components. First, $-d_{t}\psi_{E}$ quantifies the rate of change of free energy due to the external work needed for driving $E$ through control parameter change. Second, $-d_{t}D^{KL}_{{E}}(\mathcal{P}[\{\rho\}])$ quantifies the EPR due to the relaxation of $E$. Third, $\langle\dot{\Sigma}^{nr}\rangle$ is the non-reciprocal EPR whose case-specific simplifications for the $\mathcal{PT}$ preserving-breaking phases have been discussed before. Fourth, $\langle\dot{\Sigma}^{ch}\rangle+\langle\dot{\Sigma}^{sp}\rangle$ quantifies the EPR due to the non-conservative forces along the transition, namely the self-propulsion and enzymatic driving for the diffusive and reactive systems respectively.
@@
The orthogonal decomposition for dissipation functions that are quadratic in the driving affinity, and thus a linear relation between the current and affinity has been proven and rigorously studied [139, 140, 87, 141]. In contrast, non-linear relation between $J_{\gamma\gamma^{\prime}}^{ch}$ and $F_{\gamma\gamma^{\prime}}^{ch}$ (or $\vec{J}_{i}^{sp}$ and $\vec{F}_{i}^{sp}$) leads to non-quadratic dissipation functions in eq. 41. It gives exact and tighter bounds on EPR. A more rigorous proof of the orthogonal decomposition for non-quadratic dissipation functions has been derived in Ref. [142, 143, 144, 145, 146, 147] and it’s implications has been studied in Ref. [148, 149, 150, 151, 152, 153, 154, 155]. The proof relies on the dynamical large deviation approach [103, 156, 157]. In Ref.[86], we show that the large deviation functional of the non-reciprocal systems is the same as the one used in Ref.[142, 143, 144, 145, 146, 147]. This justifies rigorously orthogonal decomposition for non-reciprocal systems. In contrast to previous works, the novelty of our approach lies in the proposal of orthogonal decomposition in state-space ($\mu_{i}$ and $\epsilon_{i}^{\#}$) and transition-space ($F_{\gamma\gamma^{\prime}}^{ch},\vec{F}_{i}^{sp}$ and $f_{\gamma\gamma^{\prime}}^{ch},\vec{f}_{i}^{sp}$) for both microscopic and macroscopic systems. The anti-symmetric part in state-space and transition-space gives $\langle\dot{\Sigma}^{nr}\rangle$ and $\langle\dot{\Sigma}^{ch}\rangle,\langle\dot{\Sigma}^{sp}\rangle$ respectively. The symmetric part gives the rate of change of the macroscopic stochastic Massieu potential $-d_{t}\psi_{E}-d_{t}D^{KL}_{{E}}(\mathcal{P}[\{\rho\}])$. Our formulation reveals similarities and differences in the underlying thermodynamic geometrical structure between the nonreciprocal and reciprocal systems.
@@
Importantly, the macroscopic self-propulsion EPR with $l=1$, $\langle\dot{\Sigma}_{i}^{sp}\rangle=2d_{i}^{\parallel}F_{i}^{sp}\tanh{(|F_{i}^{sp}|/2)}$ is bounded below by the continuum space macroscopic self-propulsion EPR $\dot{\Sigma}_{i}^{sp}=\tilde{d}_{i}^{\parallel}\left(f_{i}^{sp}\right)^{2}$ in the continuum limit $l\to 0$ [86]. It highlights the importance of the observation length-scale for the coarse-grained macroscopic description. The macroscopic continuum description oversimplifies the coarse-grained description beyond the inherent diffusive length-scale of the system. Thus, it results in the underestimation of the thermodynamic dissipation due to non-conservative self-propulsion forces [69, 70, 71]. This is realized by the quadratic dissipation function in the continuum limit. The correct identification of the discreteness/finiteness of the diffusive length-scale allows the restoration of the exact microscopic dissipation at the macroscale.
D Temporal cross-correlations between macrostates and relaxation EP

### v3

@@ A Macroscopic Thermodynamics
The total mean macroscopic EPR is $\langle\dot{\Sigma}\rangle=\langle\dot{\Sigma}^{\mathcal{D}}\rangle+\langle\dot{\Sigma}^{\mathcal{R}}\rangle+d_{t}S^{gb}$. It is decomposed into the reactive mean macroscopic bulk EPR ($\langle\dot{\Sigma}^{\mathcal{R}}\rangle$), the diffusive mean macroscopic bulk EPR ($\langle\dot{\Sigma}^{\mathcal{D}}\rangle$), and the boundary term, i.e., the Gibbs EPR ($d_{t}S^{gb}$):
@@
Here, $\mathbb{D}[\{\rho\}]$ denotes the path integral over the macrostate space, and $\mathcal{P}[\{\rho\}]$ is the probability distribution for the macrostate. The mean EPR in eq. 34 has the form of a transition current multiplied by the corresponding transition affinity. The transition affinity is obtained using the macroscopic local detailed balance (LDB) condition in eq. 28. The macroscopic Gibbs entropy is defined as $S^{gb}=-\int\mathbb{D}[\{\rho\}]\mathcal{P}[\{\rho\}]\ln{(\mathcal{P}[\{\rho\}])}$, which is related to the macroscopic state entropy $S^{state}$, defined as [18, 20]:
@@ B Conservative and non-conservative decomposition of the total macroscopic EPR
$\langle\dot{\Sigma}\rangle$ is decomposed into four contributions, analogous to eq. 13:
@@
Equation 36 represents the macroscopic second law of thermodynamics. The term $d_{t}\langle E\rangle+d_{t}S^{gb}$ corresponds to the conservative EPR $\langle\dot{\Sigma}^{c}\rangle$, so $\Sigma^{c}$ depends only on the initial and final states. The exact expressions for $\langle\dot{\Sigma}^{ch}\rangle$ and $\langle\dot{\Sigma}^{sp}\rangle$ are:
@@
The non-reciprocal EPR $\langle\dot{\Sigma}^{nr}\rangle$ consists of reactive $\langle\dot{\Sigma}_{\mathcal{R}}^{nr}\rangle=\int_{\mathcal{V}}\sum_{\{\gamma\gamma^{\prime}\}}\langle J_{\gamma\gamma^{\prime}}\left(F^{nr}_{\gamma^{\prime}}-F^{nr}_{\gamma}\right)\rangle$ and diffusive contributions $\langle\dot{\Sigma}^{nr}_{\mathcal{D}}\rangle=-\int_{\mathcal{V}}\sum_{i}\langle\vec{J}_{i}^{\vec{\mathcal{D}}}\cdot\nabla F^{nr}_{i}\rangle$, such that $\dot{\Sigma}^{nr}=\dot{\Sigma}_{\mathcal{R}}^{nr}+\dot{\Sigma}^{nr}_{\mathcal{D}}$. Using the EOM eq. 29 leads to $\dot{\Sigma}^{nr}=-\sum_{i}\int_{\mathcal{V}}F_{i}^{nr}\>\partial_{t}\rho_{i}$. With $V_{ij}^{nr}=-V_{ji}^{nr}$, this further simplifies to:
@@
Here, $\omega_{ij}=\rho_{i}\partial_{t}\rho_{j}-\rho_{j}\partial_{t}\rho_{i}$ defines the macroscopic vorticity between $\rho_{i}$ and $\rho_{j}$. $\langle{\Sigma}^{nr}\rangle$ corresponds to the sustained vorticity currents between the macrostates generated by the macroscopic non-reciprocal forces. Importantly, the vorticity currents are defined in the macrostate space, in contrast to the dissipative currents defined in the transition space. Moreover, vorticity currents are defined using two macrostates, unlike chemical reaction networks, which require a transition cycle of three states to formulate a dissipative current [1]. From a fundamental perspective, this reveals a sophisticated mechanism of non-equilibriumness in non-reciprocal systems, with no equivalent counterparts in the known literature.
C Non-reciprocal phase transitions are dynamical phase transitions
The dynamics of non-reciprocal systems exhibit dynamic phases due to $\mathcal{PT}$ symmetry breaking [54], for instance, traveling waves and temporal oscillations. Here, we focus on non-reciprocal phase transitions that exhibit a transition from a static to a dynamic phase. A static phase is defined as the $\mathcal{PT}$ symmetry-preserving phase; in contrast, the $\mathcal{PT}$ symmetry is broken in the dynamic phase. We consider a wave solution $\rho_{i}=\rho_{i}^{avg}+\phi_{i}^{ow}\cos{(\theta_{i}^{ow}-v^{ow}t)}+\phi_{i}^{tw}\cos{(x-v^{tw}t+\theta_{i}^{tw})}$. Here, $\rho_{i}^{avg}$ is the time-integrated and space-integrated average value of $\rho_{i}$, and $w=\{tw,ow\}$ denotes the traveling wave and the temporal oscillating wave, respectively. The $\phi_{i}^{ow}$ and $\phi_{i}^{tw}$ characterize the wave amplitudes. Similarly, $v^{ow}$ and $v^{tw}$ are the wave velocities. Integrating eq. 38 over the space and oscillation time period $\tau^{ow}$, the non-reciprocal EP for the system reads:
@@ C Non-reciprocal phase transitions are dynamical phase transitions
Equation 39 quantifies the non-reciprocal EP density over the total volume and time period. It has the physical meaning of the kinetic energy for a wave, where the mass is given by $v^{w},V_{ij}^{nr}$ and $\sin{(\theta_{j}^{w}-\theta_{i}^{w})}$, and the velocity is given by the wave amplitudes $\phi_{i}^{w}$. For $V_{ij}^{nr}\geq 0$, it physically implies that $\rho_{i}$ repels $\rho_{j}$ while $\rho_{j}$ is attracted to $\rho_{i}$, thus $\theta_{j}^{tw}\geq\theta_{i}^{tw}$ and $\theta_{j}^{ow}\geq\theta_{i}^{ow}$, which further implies $\langle{\Sigma}^{nr}\rangle\geq 0$. Importantly, the out-of-phase state $\theta_{j}^{tw}-\theta_{i}^{tw}=\pi/2$ maximizes $\langle{\Sigma}^{nr}\rangle$, while the in-phase state $\theta_{j}^{tw}=\theta_{i}^{tw}$ minimizes $\langle{\Sigma}^{nr}\rangle$. The in-phase to out-of-phase steady-state transition has been a key motif of non-reciprocal systems [54]. Hence, it signifies that the transition from the in-phase to out-of-phase steady-state is equivalent to switching from the minimum to the maximum non-reciprocal EPR for steady-state selection. Moreover, $\langle\Sigma^{nr}\rangle$ for out-of-phase oscillation scales with the observation time $\tau$. This signifies a constant thermodynamic dissipation cost required to sustain the vorticity currents between the macrostates. Physically, it connects the dynamical phase transition to an analogy with equilibrium phase transitions, where the EPR for the non-equilibrium phase transition is analogous to the free energy for the equilibrium phase transition [141]. It is characterized by different scaling regimes for $\langle\Sigma^{nr}\rangle$. The dissipative nature of the vorticity currents is realized only in the dynamic phase. Importantly, non-reciprocal phase transitions are equivalent to dynamical phase transitions and are characterized by non-analytical behavior of $\langle\Sigma^{nr}\rangle$.
The scaling of different contributions of $\Sigma$ with the observation time $\tau$ is summarized in table 3. The self-propulsion and chemical driving EP follow scaling $O(\tau)$ for dissipative currents [89, 90]. However, a discontinuity (or a kink) in the self-propulsion or chemically driven EPR can be observed and is associated with a dynamical phase transition [141, 142, 143, 144]. The non-reciprocal phase transitions are characterized by different scaling regimes for $\Sigma^{nr}$, which correspond to sustaining temporal or spatial vorticity currents.
@@ D Orthogonal decomposition of the EPR
Using the definition in eq. 40, $d_{t}\langle{E}\rangle+d_{t}S^{gb}$ in eq. 36 is further decomposed into the orthogonal form [101]:
@@
$\psi_{E}$ is the macroscopic free energy defined as $\psi_{E}=-\log{(\mathcal{Z}^{E})}$. Here, $\mathcal{Z}^{E}=\int\mathbb{D}[\{\rho\}]e^{-E}$ is the partition function, and the macroscopic Boltzmann distribution is ${\mathcal{P}^{E}[\{\rho\}]}=e^{-E\>+\Psi^{E}}$. It holds that $\langle S^{state}_{E}\rangle=-D^{KL}_{{E}}(\mathcal{P}[\{\rho\}])$. Using the symmetry of the non-conservative external driving forces $F_{\gamma\gamma^{\prime}}$ and $\vec{F}_{i}^{sp}$, namely the orthogonal decomposition of the transition affinities, eq. 37 is reduced to:
@@
Here, $\vec{J}_{i}^{sp}=2d_{i}e^{\mu_{i}}\sinh{(\vec{F}_{i}^{sp}/2)}=T_{i}^{\perp}\sinh{(\vec{F}_{i}^{sp}/2)}$ is the anti-symmetric part of $\vec{J}_{i}^{\vec{\mathcal{D}}}$ under the adjoint transformation $\vec{F}_{i}^{sp}\to-\vec{F}_{i}^{sp}$. Similarly, $J_{\gamma\gamma^{\prime}}^{ch}=d_{\gamma\gamma^{\prime}}\left(e^{\mu_{\gamma}}+e^{\mu_{\gamma^{\prime}}}\right)\sinh{\left({F_{\gamma\gamma^{\prime}}^{ch}}/{2}\right)}=T^{\perp}_{\gamma\gamma^{\prime}}\sinh{\left({F_{\gamma\gamma^{\prime}}^{ch}}/{2}\right)}$ is the anti-symmetric part of $J_{\gamma\gamma^{\prime}}$ under the adjoint transformation $F_{\gamma\gamma^{\prime}}^{ch}\to-F_{\gamma\gamma^{\prime}}^{ch}$. Importantly, $T_{i}^{\perp}$ and $T^{\perp}_{\gamma\gamma^{\prime}}$ are the scaled (by $\Omega$) variances of currents (eq. 32) in the direction orthogonal to the external driving, which are obtained by plugging in $\vec{F}_{i}^{sp}=0$ and $F_{\gamma\gamma^{\prime}}^{ch}=0$, respectively. Thus, in eq. 42, the non-conservative EPR depends on a non-linear function of the external driving force and the current variance in the direction orthogonal to the driving. Importantly, $\langle\dot{\Sigma}^{ch}\rangle$ and $\langle\dot{\Sigma}^{sp}\rangle$ depend on $e^{\mu_{i}}$ and are proportional to the macrostate mobility.
@@
Equation 41 formulates the orthogonal decomposition of $\langle\dot{\Sigma}\rangle$. It decomposes $\langle\dot{\Sigma}\rangle$ into four linearly independent components. First, $d_{t}\psi_{E}$ quantifies the rate of change of the free energy due to the external work needed to drive $E$ through the change in control parameters. Second, $d_{t}D^{KL}_{{E}}(\mathcal{P}[\{\rho\}])$ quantifies the EPR due to the relaxation of $E$. Third, $\langle\dot{\Sigma}^{nr}\rangle$ is the non-reciprocal EPR, whose case-specific simplifications for the $\mathcal{PT}$ preserving-breaking phases have been discussed before. Fourth, $\langle\dot{\Sigma}^{ch}\rangle+\langle\dot{\Sigma}^{sp}\rangle$ quantifies the EPR due to the non-conservative forces along the transitions, namely the self-propulsion and enzymatic driving for the diffusive and reactive systems, respectively.
@@
The orthogonal decomposition for dissipation functions that are quadratic in the driving affinity, and thus a linear relationship between the current and the affinity, has been proven and rigorously studied [145, 146, 88, 147]. In contrast, the non-linear relation between $J_{\gamma\gamma^{\prime}}^{ch}$ and $F_{\gamma\gamma^{\prime}}^{ch}$ (or $\vec{J}_{i}^{sp}$ and $\vec{F}_{i}^{sp}$) leads to non-quadratic dissipation functions in eq. 42. It gives exact and tighter bounds on the EPR. A more rigorous proof of the orthogonal decomposition for non-quadratic dissipation functions has been derived in Refs. [148, 149, 150, 151, 152, 153], and its implications have been studied in Refs. [154, 155, 156, 157, 158, 159, 160, 161]. The proof relies on the dynamical large deviation approach [109, 162, 163]. In Ref.[87], we show that the large deviation functional of the non-reciprocal systems is the same as the one used in Refs. [148, 149, 150, 151, 152, 153]. This rigorously justifies the orthogonal decomposition for non-reciprocal systems. In contrast to previous works, the novelty of our approach lies in the proposal of orthogonal decomposition in state-space ($\mu_{i}$ and $\epsilon_{i}^{\#}$) and transition-space ($F_{\gamma\gamma^{\prime}}^{ch},\vec{F}_{i}^{sp}$ and $f_{\gamma\gamma^{\prime}}^{ch},\vec{f}_{i}^{sp}$) for both microscopic and macroscopic systems. The anti-symmetric part in state-space and transition-space gives $\langle\dot{\Sigma}^{nr}\rangle$ and $\langle\dot{\Sigma}^{ch}\rangle,\langle\dot{\Sigma}^{sp}\rangle$, respectively. The symmetric part gives the rate of change of the macroscopic stochastic Massieu potential $d_{t}\psi_{E}-d_{t}D^{KL}_{{E}}(\mathcal{P}[\{\rho\}])$. Our formulation reveals similarities and differences in the underlying thermodynamic geometrical structure between the non-reciprocal and reciprocal systems. Importantly, the orthogonal decomposition of the thermodynamic cost in eq. 41 is completely determined by physically measurable quantities, in particular, external driving forces and the corresponding current means and variances.
@@
Importantly, the macroscopic self-propulsion EPR with $l=1$, $\langle\dot{\Sigma}_{i}^{sp}\rangle=2d_{i}^{\parallel}F_{i}^{sp}\tanh{(|F_{i}^{sp}|/2)}$, is bounded below by the continuum-space macroscopic self-propulsion EPR $\dot{\Sigma}_{i}^{sp}=\tilde{d}_{i}^{\parallel}\left(f_{i}^{sp}\right)^{2}$ in the continuum limit $l\to 0$ [87]. This highlights the importance of the observation length scale for the coarse-grained macroscopic description. The macroscopic continuum description oversimplifies the coarse-grained description beyond the inherent diffusive length scale of the system, resulting in an underestimation of thermodynamic dissipation due to non-conservative self-propulsion forces [69, 70, 71]. This underestimation arises from the quadratic dissipation function in the continuum limit. Correctly identifying the discreteness/finiteness of the diffusive length scale restores the exact microscopic dissipation at the macroscale.
E Temporal cross-correlations between macrostates and relaxation EP